OpenStudy (anonymous):

The ratio of the areas of two similar polygons is 225:25. What is the ratio of their corresponding sides?

OpenStudy (saifoo.khan):

\[\frac{A_1}{A_2} = \frac{(L_1)^2}{(L_2)^2}\]

OpenStudy (anonymous):

What? Shouldnt it be something like 2:1 or similar?

OpenStudy (saifoo.khan):

brb

OpenStudy (anonymous):

Alright.

hero (hero):

Ratio of areas = a^2: b^2 Ratio of sides = a:b In this case... Ratio of areas = (15)^2: (5)^2 Ratio of sides = 15:5 simple

OpenStudy (anonymous):

Ahh I see now. Thanks buddy.

hero (hero):

By the way, Ratio of volume is a^3:b^3 ...Just in case you need that. You will need it at some point

OpenStudy (anonymous):

Thanks, I wrote it down.